A sample of chromium contains four isotopes and has a relative atomic mass of 52.09
s the mass number and the percentage abundance of three of these isotopes.
Mass number 52 53 54
Abundance (%) 82.8, 10.9, 2.7
Determine the percentage abundance of the fourth isotope. Show by calculation that the mass number of this isotope is
1
Expert's answer
2020-07-07T14:29:57-0400

Solution.

w(xCr)=1-w(52Cr)-w(53Cr)-w(54Cr)=1-0.828-0.109-0.027=0.036 (3.6 %)

Ar(Cr)=Ar(52Cr)*w(52Cr)+Ar(53Cr)*w(53Cr)+Ar(54Cr)*w(54Cr)+Ar(xCr)*w(xCr)

52.09=52*0.828+53*0.109+54*0.027+Ar(xCr)*0.036

Ar(xCr)=50

Answer: the mass number of this isotope is 50.